The height of students at a university taking introduction to probability are nomally distributed with mean height of 50 cm and standard deviation 10 cm. Find the probability that the height of a randomly selected standard is



(a) shorter than Wye



(b) within 5 cm of the mean

1
Expert's answer
2022-01-12T18:30:14-0500

Let "X=" height of student, "X\sim N(\mu, \sigma^2)."

Given "\mu=50cm, \sigma=10cm"

(a)


"P(X<W_{ye})=P(Z<\dfrac{W_{ye}-50}{10})"

(b)


"P(|X-\mu|<2.5)=P(47.5<X<52.5)"

"=P(X<52.5)-P(X\leq4705)"

"=P(Z<\dfrac{52.5-50}{10})-P(Z\leq\dfrac{47.5-50}{10})"

"=P(Z<0.25)-P(Z\leq-0.25)"

"\approx0.598706-0.401294"

"\approx0.1974"